Tilting modules and Gorenstein rings

Lidia Angeleri Hügel, Dolors Herbera and Jan Trlifaj ^{∗}

In his pioneering work [30], Miyashita extended classical tilting theory to the setting of finitely presented modules of finite projective dimension over an arbitrary ring. More recently, tilting theory has been generalized to arbitrary modules of finite projective dimension [2], [3], [6], et al. Encouraged by this developement, in the present paper, we apply infinite dimensional tilting theory to the setting of modules over Iwanaga-Gorenstein rings.

We start with pointing out the central role played by resolving subcategories of modR for an arbitrary ring R. In Theorem 2.2 we prove that these subcategories correspond bijectively to tilting classes of finite type in ModR, as well as to cotilting classes of cofinite type in RMod. This generalizes a classical result of Auslander-Reiten [9] characterizing cotilting classes in modR over an artin algebra R. Moreover, it provides for a bijection between certain cotilting left modules and tilting right modules which extends the duality between finitely generated cotilting and tilting modules over artin algebras.

We then consider Iwanaga-Gorenstein rings, that is (not necessarily commutative) two-sided noetherian rings of finite self-injective dimension on both sides. We show that the first finitistic dimension conjecture holds true for any Iwanaga-Gorenstein ring (Theorem 3.2). We prove that Gorenstein injectivity, and flatness, can be tested by finitely generated modules of finite projective dimension (Corollary 3.5). We also show that Iwanaga-Gorenstein rings are characterized among noetherian rings by the property that Gorenstein injective modules on either side form a tilting class (Theorem 3.4). For an artin algebra R, we obtain that R is Iwanaga-Gorenstein if and only if the class of Gorenstein projective right modules is induced by a finitely generated cotilting module (Corollary 3.8).

∗This work was started while the first author was a Ramon y Cajal Fellow at the Universitat Autònoma de Barcelona. The research of the second author was funded by the DGI (Spain) and FEDER through Project BFM2002-01390, and by the Comissionat per Universitats i Recerca of the Generalitat de Catalunya. Third author supported by grant SAB2001-0092 of Secretaria de Estado de Education y Universidades MECD at CRM IEC Barcelona, and by MSM 113200007.

In the final section, we provide for an explicit construction of tilting and cotilting modules over any commutative 1-Gorenstein ring R (Example 4.1). By [32], it is consistent with ZFC that these are the only tilting and cotilting modules up to equivalence in case R is a small Dedekind domain.

1 Preliminaries

Let R be a ring. Denote by ModR the category of all (right R-) modules, and by modR the subcategory of all modules possessing a projective resolution consisting of finitely generated modules. The elements of modR are sometimes referred to as the modules of type FP_{∞}, cf. [12, VIII, §4]; if R is right coherent then these are exactly the finitely presented modules.

We denote by P and I the class of all modules of finite projective and injective dimension, respectively. For n < ∞, we denote by P_{n} (I_{n}, F_{n}) the class of all modules of projective (injective, flat) dimension ≤ n.

Finally, we put P^{<∞} = P∩ modR, P_{n}^{<∞} = P_{n}∩ modR, and I_{n}^{<∞} = I_{n} ∩ modR.

\underline{A. Resolving subcategories.} Let C be a class of finitely generated modules. Then C is resolving provided that (R1) C contains all finitely generated projective modules, (R2) C is closed under direct summands and extensions, and (R3) X ∈ C whenever there is an exact sequence 0 ⟶ X ⟶ Y ⟶ Z ⟶ 0 with Y, Z ∈ C.

Notice that every resolving class C is contained in modR. Moreover, the property of being resolving can be tested in a weaker form:

Lemma 1.1 Let R be a ring.

1. Let C ⊆ modR. Then C is resolving iff C satisfies conditions (R1), (R2) and (R3’) X ∈ C whenever there is an exact sequence 0 ⟶ X ⟶ P ⟶ Y ⟶ 0 such that P is finitely generated projective and Y ∈ C.

2. The class modR is resolving.

3. Assume C ⊆ P_{1}^{<∞}. Then C is resolving iff C satisfies conditions (R1) and (R2).

Proof : 1. Clearly, any resolving class C satisfies (R3’). Conversely, assume C satisfies (R1), (R2) and (R3’). In order to prove (R3), consider an exact sequence 0 ⟶ X ⟶ Y ⟶ Z ⟶ 0 with Y, Z ∈ C. By assumption, there is an exact sequence 0 ⟶ K ⟶ P ⟶ Z ⟶ 0 with P finitely generated projective and K ∈ C. Consider the pull-back diagram

0  0
↓  ↓
K  K
↓  ↓
0 ⟶ X ⟶ L ⟶ P ⟶ 0
∥  ↓  ↓
0 ⟶ X ⟶ Y ⟶ Z ⟶ 0
↓  ↓
0  0

Since C is closed under extensions, the left column gives L ∈ C. Since P is projective, the upper row splits, X is a direct summand of L, and hence X ∈ C. This proves condition (R3).

2. Condition (R1) is obvious. Schanuel’s lemma gives closure of modR under direct summands, and the horseshoe lemma implies closure under extensions - so (R2) holds. Schanuel’s lemma then gives (R3’), and part 1. applies.

3. If C ⊆ modR ∩ P_{1}, then (R3’) follows from (R1). □

\underline{B. Orthogonal classes.} For a class C ⊆ ModR, we define C^{⊥_{1}} = KerExt_{R}^{1} (C, −), ⊥_{1}C = KerExt_{R}^{1} (−, C), C^{⊥} = ⋂_{i≥1} Ker Ext_{R}^{i}(C, −), and ^{⊥}C = ⋂_{i≥1} Ker Ext_{R}^{i}(−, C). Similarly, we define the classes C^{⊺_{1}}, ^{⊺_{1}}C, C^{⊺}, and ^{⊺}C, replacing Ext by Tor.

Observe that ^{⊥}C and ^{⊺}C satisfy conditions (R1), (R2), (R3) above. So, it follows from Lemma 1.1.2 that ^{⊥}C ∩ modR and ^{⊺}C ∩ modR are resolving.

We collect here some further properties which we will often use in the sequel.

Lemma 1.2 If C ⊆ modR, then C^{⊥} is closed under direct limits, and C^{⊺} is closed under direct products.

Proof : See [19, 10.2.4 and 3.2.26].□

Theorem 1.3 Let C ⊆ modR be a resolving class, and denote by \underrightarrow{lim} C the class of all modules that are direct limits of an inductive system of modules from C. Then the following hold true:

(1) C^{⊥} = C^{⊥_{1}}, and C^{⊺} = C^{⊺_{1}}.

(2) ^{⊥}(C^{⊥}) ⊆ \underrightarrow{lim} C = ^{⊺}(C^{⊺}).

(3) C = ^{⊥}(C^{⊥}) ∩ modR = \underrightarrow{lim} C ∩ modR = ^{⊺}(C^{⊺}) ∩ modR.

Proof : see [2, 1.2], [7, 2.3] and [27].□

We shall denote by S^{∗} = Hom_{ℤ}(S, ℚ/ℤ) the dual of the module S. As we recall in the next lemma this definition of dual module has useful homological properties. For particular rings, other notions with the same properties can be considered: the k-dual module in case R is an Artin k-algebra, and the W -dual module where W is any injective cogenerator for ModR in case R is commutative.

Lemma 1.4 Let R be a ring, A, D ∈ ModR, B ∈ RMod, and C ∈ modR. Let i ≥ 1. Then

1. Ext_{R}^{i}(A, B^{∗}) ≅ (Tor_{i}^{R}(A, B))^{∗} and

2. Tor_{i}^{R}(C, D^{∗}) ≅ (Ext_{R}^{i}(C, D))^{∗}.

Proof : See e.g. [19, 3.2.1 and 3.2.13]. □

\underline{C. Cotorsion pairs.} A pair of classes of modules ℭ = (A, B) is a cotorsion pair provided that A = ^{⊥_{1}}B and B = A^{⊥_{1}}. A cotorsion pair ℭ = (A, B) is complete provided that for each module M there is an exact sequence 0 → M → B → A → 0 with A ∈ A and B ∈ B (or, equivalently, provided that for each module M′ there is an exact sequence 0 → B′ → A′ → M′ → 0 with A′ ∈ A and B′ ∈ B).

For example, (P_{n}, P_{n}^{⊥}), (^{⊥}I_{n}, I_{n}) and (F_{n}, F_{n}^{⊥}) are complete cotorsion pairs for any ring R, cf. [19, §7.4], [31] and [17].

A cotorsion pair ℭ = (A, B) in ModR is of finite type provided there exists a class S ⊆ modR such that B = S^{⊥}. The class B is then definable, that is, it is closed under direct products, direct limits, and pure submodules.

Dually, a cotorsion pair ℭ = (A, B) in RMod is of cofinite type provided there is a class S ⊆ modR such that A = S^{⊺}. Equivalently, A = ^{⊥}S^{∗} where S^{∗} = {S^{∗} | S ∈ S}.

2 Tilting modules of finite type

Let n < ∞. A module T is n-tilting provided (T1) T ∈ P_{n}, (T2) Ext_{R}^{i}(T, T^{(I)}) = 0 for each i ≥ 1 and all sets I, and (T3) there exist r ≥ 0 and a long exact sequence 0 ⟶ R ⟶ T_{0} ⟶ ⋯ ⟶ T_{r} ⟶ 0 such that T_{i} ∈ AddT for each 0 ≤ i ≤ r.

A class of modules T is n-tilting provided there is an n-tilting module such that T = T^{⊥}. In this case, ℭ = (^{⊥}T , T ) is a complete cotorsion pair, cf. [16]. If ℭ is of finite type, then T and T are called n-tilting of finite type.

Clearly, T is n-tilting of finite type in case T ∈ modR. However, there are examples of tilting classes of finite type not induced by any T ∈ modR: if R is an artin algebra of finitistic dimension n < ∞ such that P^{<∞} is not contravariantly finite, then P^{<∞}^{⊥} is an n-tilting class of finite type, but there is no n-tilting module T ∈ modR such that T^{⊥} = P^{<∞}^{⊥}, cf. [6, 25]. Another instance of this fenomenon occurs in Example 4.1 below. In fact, no example appears to be known of an n-tilting class which is not of finite type, or even not definable.

Dually, a module C is n-cotilting provided that (C1) C ∈ I_{n}, (C2) Ext_{R}^{i}(C^{I}, C) = 0 for each i ≥ 1 and all sets I, and (C3) there exists r ≥ 0 and a long exact sequence 0 ⟶ C_{r} ⟶ ⋯ ⟶ C_{0} ⟶ W ⟶ 0 where W is an injective cogenerator for ModR and C_{i} ∈ ProdC for each 0 ≤ i ≤ r.

A class of modules F is n-cotilting provided there is an n-cotilting module such that F = ^{⊥}C. In this case, ℭ = (F, F^{⊥}) is a complete cotorsion pair, cf. [2]. If ℭ is of cofinite type, then F and C are called n-cotilting of cofinite type.

Tilting classes and cotilting classes can be characterized as follows.

Theorem 2.1 [2, 4.2 and 4.3]

(1) A class of modules T is n-tilting if and only if

(a) (^{⊥}T , T ) is a complete cotorsion pair,

(b) T is closed under arbitrary direct sums, and X ∈ T whenever there is an exact sequence 0 ⟶ Z ⟶ Y ⟶ X ⟶ 0 with Y, Z ∈ T ,

(c) ^{⊥}T ⊆ P_{n}.

(2) A class of modules F is n-cotilting if and only if (a) (F, F^{⊥}) is a complete cotorsion pair,

(b) F is closed under arbitrary direct products, and X ∈ F whenever there is an exact sequence 0 ⟶ X ⟶ Y ⟶ Z ⟶ 0 with Y, Z ∈ F,

(c) F^{⊥} ⊆ I_{n}.

In fact, the Theorem above was inspired by a previous result due to Auslander and Reiten [9, Theorem 5.5] which characterizes cotilting classes in modR – in case R is an artin algebra – as the contravariantly finite resolving subcategories F of modR satisfying F^{⊥} ∩ modR ⊆ I_{n}^{<∞}.

The characterization gets simpler if we consider (co)tilting classes of (co)finite type in the category of arbitrary modules. In fact, we will now show that there is a bijective correspondence between n-tilting classes of finite type in ModR and the resolving subclasses in P_{n}^{<∞}, and there is a similar correspondence for n-cotilting classes of cofinite type in RMod. The point is that there are more approximations available in ModR - for instance, any cotorsion pair cogenerated by a set of modules is complete, cf.[16].

Theorem 2.2 Let R be a ring and n < ∞.

1. The n-tilting classes of finite type in ModR correspond bijectively to the resolving subclasses of P_{n}^{<∞}. The correspondence is given by the mutually inverse assignments α : T ↦ ^{⊥}T ∩ modR and β : C ↦ C^{⊥}.

2. The n-cotilting classes of cofinite type in RMod correspond bijectively to the resolving subclasses of P_{n}^{<∞}. The correspondence is given by the mutually inverse assignments γ : F ↦ ^{⊺}F∩ modR and δ : C ↦ C^{⊺}.

3. Let S ⊆ P_{n}^{<∞}. Then T = S^{⊥} is an n-tilting class of finite type, and F = S^{⊺} is the corresponding n-cotilting class of cofinite type, that is, F = δα(T ).

Proof : 1. For any n-tilting class T we have ^{⊥}T ⊆ P_{n} by Theorem 2.1, so α(T ) is a resolving subclass of P_{n}^{<∞}. Next, let C be a resolving subclass of P_{n}^{<∞}. Then C^{⊥} is closed under direct sums by Lemma 1.2, and (^{⊥}(C^{⊥}), C^{⊥}) is a complete cotorsion pair by [16]. Moreover, C^{⊥} is closed under cokernels of monomorphisms, and ^{⊥}(C^{⊥}) ⊆ P_{n} by [2, Lemma 2.2(a)]. So β(C) = C^{⊥} is an n-tilting class (of finite type) by Theorem 2.1.

Now, let T be an n-tilting class of finite type. Take S ⊆ modR such that S^{⊥} = T . Then S ⊆ α(T ), so βα(T ) ⊆ T . On the other hand, α(T ) ⊆ ^{⊥}T , so T ⊆ (^{⊥}T )^{⊥} ⊆ βα(T ), and βα =id. Conversely, if C is a resolving subclass of P_{n}^{<∞}, then Theorem 1.3.3 proves that αβ =id.

2. Let F be an n-cotilting class of cofinite type, and consider S ⊆ modR such that F = S^{⊺}(= ^{⊥}S^{∗}). Then S^{∗} ⊆ F^{⊥} ⊆ I_{n} by Theorem 2.1. So S ⊆ F_{n}, and γ(F) ⊆ ^{⊺}(S^{⊺}) ⊆ ^{⊺}(F_{n}^{⊺}) = F_{n} by [7, 2.7]. Since γ(F) ⊆ modR, each F ∈ γ(F) has a partial projective resolution 0 ⟶ Q ⟶ P_{n−1} ⟶ ⋯ ⟶ P_{0} ⟶ F ⟶ 0 where P_{i} (i < n) are finitely generated projective, and Q is (finitely presented and) flat. Then Q is projective, and F ∈ P_{n}. This proves that γ(F) is a resolving subclass of P_{n}^{<∞}.

On the other hand, if C is a resolving subclass of P_{n}^{<∞}, then C^{⊺} = ^{⊥}C^{∗} gives rise to a complete cotorsion pair (C^{⊺}, (C^{⊺})^{⊥}) by [17, Corollary 10]. Moreover, C^{∗} ⊆ I_{n}, so (C^{⊺})^{⊥} = (^{⊥}C^{∗})^{⊥} ⊆ I_{n} by [2, Lemma 2.2(b)]. Finally, C^{⊺} satisfies (R3), and it is closed under direct products by Lemma 1.2. It now follows from Theorem 2.1 that δ(C) is an n-cotilting class (of cofinite type).

The proof for γδ=id and δγ=id is analogous to that in 1.

3. Let C be the smallest resolving subclass of P_{n}^{<∞} containing S. Then T = S^{⊥} = C^{⊥}, and so F = S^{⊺} = C^{⊺} by Lemma 1.4.1. By part 1. and 2. we know that T = β(C) is a tilting class and F = δ(C) is a cotilting class. Moreover, F = δαβ(C) = δα(T ). □

Besides the correspondence for the classes of modules, there is also a correspondence for the individual modules:

Proposition 2.3 Let R be a ring and n < ∞. Let T be an n-tilting module of finite type in ModR. Let T = T^{⊥}, and let F = δα(T ) be the corresponding n-cotilting class.

Then C = T^{∗} is an n-cotilting module of cofinite type in RMod with ^{⊥}C = F = T^{⊺}. Moreover, M ∈ T iff M^{∗} ∈ F, for each M ∈ ModR.

Proof : Applying Hom_{ℤ}(−, ℚ/ℤ) to the long exact sequences in (T1) and (T3) for T, we obtain conditions (C1) and (C3) for C.

By Theorem 2.2.1, we further know that S = α(T ) ⊆ P_{n}^{<∞} is resolving and S^{⊥} = T = T^{⊥}. In particular, S^{⊥} ∩ (RMod)^{∗} = T^{⊥} ∩ (RMod)^{∗}, and hence S^{⊺} = T^{⊺} by Lemma 1.4.1.

Consider M ∈ModR. By (the left-sided version of) Lemma 1.4.1, for each i > 0, Ext_{R}^{i}(M^{∗}, C) = 0 iff Tor_{i}^{R}(T, M^{∗}) = 0. It follows that M^{∗} ∈ ^{⊥}C iff M^{∗} ∈ S^{⊺} iff M ∈ S^{⊥} = T , by Lemma 1.4.2. Taking M = T^{(I)} for an arbitrary set I, we see that condition (C2) for C is equivalent to (T2) for the module T.

Finally, we have F = δαβ(S) = δ(S) = S^{⊺}, while ^{⊥}C = T^{⊺} = S^{⊺}. □

Let us notice some immediate consequences of Theorem 2.2 in particular settings:

Examples 2.4 (i) Let R be a ring. Then 1-tilting (1-cotilting) classes of finite (cofinite) type correspond bijectively to subclasses of P_{1}^{<∞} satisfying conditions (R1) and (R2). This follows by Lemma 1.1.3.

(ii) Let R be a von Neumann regular ring. Then modR ⊆ P_{0}. So, a module T is tilting of finite type (cotilting of cofinite type) iff T is a projective generator (injective cogenerator). Moreover, ModR is the only tilting (cotilting) class of finite (cofinite) type.

(iii) Let R be a right semihereditary ring (a Prüfer domain, for example). Then modR ⊆ P_{1}. So, any tilting module of finite type has projective dimension ≤ 1, and any cotilting module of cofinite type has injective dimension ≤ 1.

(iv) Let R be an elementary divisor domain (a semilocal Prüfer domain, for example), see [22, III.§6]. A class T ⊆ ModR is tilting of finite type iff there is a set S ⊆ R such that T coincides with the class of all S-divisible modules (that is, of all modules M such that sM = M for all s ∈ S, or equivalently, Ext_{R}^{1}(R/sR, M) = 0 for all s ∈ S).

This is because each finitely presented module is a direct sum of cyclically presented modules [22, V.3.4]. So a resolving subclass C ⊆ modR is determined by the set S = {s ∈ R | R/sR ∈ C}, and Theorem 2.2.3 applies.

Similarly, a class F ⊆ ModR is cotilting of cofinite type iff there is a set S ⊆ R such that F coincides with the class of all S-torsion-free modules (that is, of all modules M such that sm ≠ 0 for all 0 ≠ m ∈ M and 0 ≠ s ∈ S, or equivalently, Tor_{1}^{R}(R/sR, M) = 0 for all s ∈ S). □

We remark that resolving subclasses of P_{1}^{<∞} have good approximation properties in the noetherian setting:

Proposition 2.5 Let R be a right noetherian ring. Let C be a subclass of P_{1}^{<∞} which is closed under direct summands and extensions and contains R. Then C is covariantly finite.

Proof : The class C is resolving by Lemma 1.1.3. Put D = C^{⊺}. Then ^{⊺}D = \underrightarrow{lim} C by Theorem 1.3. By assumption, D is closed under submodules, and each module in D is a direct limit of finitely presented modules in D. It follows that ^{⊺}D = ^{⊺}(D∩ modR). The latter class is closed under direct products by Lemma 1.2, so [15, 4.2] shows that C is covariantly finite. □

Tilting classes of finite type can also be characterized as certain definable classes:

Proposition 2.6 Let R be a ring.

1. Let T be a tilting class and T be a tilting module such that T^{⊥} = T . Then T is of finite type iff T is definable and T ∈ \underrightarrow{lim} (^{⊥}T ∩ modR).

2. Assume R is right noetherian. Let F be a 1-cotilting (torsion-free) class and C be a 1-cotilting module such that ^{⊥}C = F. Then F is of cofinite type iff C^{∗} ∈ \underrightarrow{lim} (^{⊺}F∩ modR).

Proof : 1. By Theorem 2.1, the class C = ^{⊥}T ∩ modR is a resolving subclass of P_{n}^{<∞} where n is the projective dimension of T. If T is of finite type, then by Theorem 2.2.1 we have T = C^{⊥}, hence T is definable. Moreover, T ∈ ^{⊥}(C^{⊥}) ⊆ \underrightarrow{lim} C by Theorem 1.3.

Conversely, put A = C^{⊥}. Then T ⊆ A, both A and T are definable, and A is tilting of finite type by Theorem 2.2.1. In order to prove that A = T , it suffices to show that A and T contain the same pure-injective modules.

Let P ∈ A be pure-injective. By assumption, there is a direct system (C_{σ} | σ ∈ I) of modules in C such that T = \underrightarrow{lim}_{σ∈I} C_{σ}, and Ext_{R}^{i}(C_{σ}, P) = 0 for all σ ∈ I and i > 0. Then Ext_{R}^{i}(T, P) ≅ \underleftarrow{lim}_{σ∈I}Ext_{R}^{i}(C_{σ}, P) = 0 (i > 0), hence P ∈ T .

2. First, we prove that ^{⊥_{1}}C = ^{⊥_{1}}C^{∗∗}. Let F ∈ modR. By Lemma 1.4, we have Ext_{R}^{1}(F, C^{∗∗}) = 0 iff Tor_{1}^{R}(F, C^{∗}) = 0 iff Ext_{R}^{1}(F, C) = 0, so ^{⊥_{1}}C∩ modR = ^{⊥_{1}}C^{∗∗}∩ modR.

By [11], C is pure-injective, so C is a direct summand in C^{∗∗}, hence ^{⊥_{1}}C^{∗∗} ⊆ ^{⊥_{1}}C. Conversely, let M ∈ ^{⊥_{1}}C. Since ^{⊥_{1}}C is a torsion-free class and R is right noethe-

rian, M = \underrightarrow{lim}_{σ∈I} C_{σ} for a directed system (C_{σ} | σ ∈ I) of finitely presented modules in ⊥_{1}C. Then C_{σ} ∈ ^{⊥_{1}}C^{∗∗} for all σ ∈ I, and Ext_{R}^{1}(M, C^{∗∗}) ≅ \underleftarrow{lim}_{σ∈I}Ext_{R}^{i}(C_{σ}, C^{∗∗}) = 0.

This proves that F = ^{⊥_{1}}C = ^{⊥_{1}}C^{∗∗} = (C^{∗})^{⊺_{1}}.

Let now C = ^{⊺}F∩ modR. If F is of cofinite type, then by Theorem 2.2.2. we have C^{⊺} = F = (C^{∗})^{⊺_{1}}, and C^{∗} ∈ ^{⊺_{1}}(C^{⊺_{1}}) = \underrightarrow{lim} C by Theorem 1.3.

Conversely, assume C^{∗} ∈ \underrightarrow{lim} C. Since \underrightarrow{lim} C = ^{⊺_{1}}(C^{⊺_{1}}), we have F = (C^{∗})^{⊺_{1}} = C^{⊺_{1}}, so F is of cofinite type. □

Since the classes P_{0}^{<∞} and P_{n}^{<∞} are resolving there always exists the least, and the largest, n-tilting class of finite type (and similarly for the n-cotilting classes). While the largest tilting class is induced by any progenerator, the least class may not be induced by any tilting module in modR (cf. [6, 4.4]). Further examples of tilting classes of finite type not induced by tilting modules in modR will be exhibited in the next two sections.

3 Iwanaga-Gorenstein rings

Recall that a ring R is said to be Gorenstein, or Iwanaga-Gorenstein [19], if R is (left and right) noetherian and both idim R_{R} and idim _{R}R are finite. It is well known that idim R_{R} and idim _{R}R then coincide [19, 9.1.8], and R is called n-Gorenstein if this number is n. Over an n-Gorenstein ring, one has P = P_{n} = I = I_{n}, see [19, 9.1.10]. In particular, the big finitisitic dimension FindimR, that is, the supremum of the projective dimensions attained on P, equals n. We are now going to see that also the little finitistic dimension, that is, the supremum of the projective dimensions attained on P^{<∞}, equals n.

To this end, we first have to recall the following result due to Auslander and Buchweitz. If X ⊆ ModR, then X̂ will denote the subcategory of all modules M_{R} admitting a finite X -resolution, that is, a long exact sequence 0 ⟶ X_{r} ⟶ ⋯ ⟶ X_{1} ⟶ X_{0} ⟶ M ⟶ 0 with r ≥ 0 and X_{i} ∈ X for i = 0, ⋯, r.

Proposition 3.1 [8, 1.8] Let R be a noetherian ring such that _{R}R has finite injective dimension, and let X = ^{⊥}R ∩ modR. Then for each X ∈ X̂, there exist short exact sequences

0 ⟶ K_{X} ⟶ M_{X} ⟶^{g_{X}} X ⟶ 0 with K_{X} ∈ P^{<∞}, and M_{X} ∈ X ,

0 ⟶ X ⟶^{f_{X}} B_{X} ⟶ C_{X} ⟶ 0 with B_{X} ∈ P^{<∞}, and C_{X} ∈ X .

Moreover, if R is a Gorenstein ring, then X̂ = modR.

Theorem 3.2 If R is an n-Gorenstein ring, then FindimR =findimR = n, and every module of finite projective dimension is a direct summand of a P^{<∞}-filtered module.

Proof : Let R be n-Gorenstein, so P = P_{n} = I = I_{n}. Let (A, B) be the cotorsion pair cogenerated by P^{<∞}. We claim that every module N in P ∩ B is injective. By the Baer Lemma, it is enough to show that Ext_{R}^{1} (R/I, N) = 0 for every right ideal I of R. Now, setting X = R/I and X = ^{⊥}R ∩ modR, we know from Proposition 3.1 that there is a short exact sequence 0 ⟶ X ⟶^{f_{X}} B_{X} ⟶ C_{X} ⟶ 0 with B_{X} ∈ P^{<∞}, and C_{X} ∈ X . This yields a long exact sequence … Ext_{R}^{1} (B_{X}, N) → Ext_{R}^{1} (R/I, N) → Ext_{R}^{2} (C_{X}, N) … where Ext_{R}^{1} (B_{X}, N) = 0 since N ∈ B, and Ext_{R}^{2} (C_{X}, N) = 0 because N ∈ P and X^{⊥} contains all projective modules and is coresolving, hence contains P. Now, since findimR ≤ n, we know from [6, Theorem 2.6] that (A, B) is a tilting cotorsion pair induced by a tilting module T, and that A ⊆ P. Then AddT = A ∩ B ⊆ P ∩ B consists of injective modules. In particular, AddT is closed under cokernels of monomorphisms. By [6, Theorem 3.2] we then know that A = P and that our statement holds true. □

The following result was first proven in [28, 4.7]. Further proofs are given in [29] and [1]. We now obtain yet another proof.

Corollary 3.3 Let R be a Gorenstein ring and let 0 → R_{R} → I_{0} → I_{1} → … I_{n} → 0 be a minimal injective resolution of R_{R}. Then every indecomposable injective right module is isomorphic to a direct summand of some I_{k} with 0 ≤ k ≤ n.

Proof : We have just seen that (P, B) is a tilting cotorsion pair whose kernel P ∩ B is the class of the injective modules. Observe that the module I = ⨁_{i=0}^{n} I_{k} is a tilting module with I^{⊥} = B, see [2, proof of 4.1]. Then AddI = P ∩ B. In particular, every indecomposable injective module belongs to AddI and thus is isomorphic to an indecomposable direct summand of some I_{k} by Azumaya’s Decomposition Theorem. □

Recall from [18] and [19] that a module M over an arbitrary ring R is said to be Gorenstein injective if it can be written as M = Ker f for some long exact sequence

… E_{1} → E_{0} → E^{0} ⟶^{f} E^{1} → …

of injective modules which stays exact under Hom_{R}(I, ) for any injective module I_{R}. The Gorenstein projective modules are defined dually. Furthermore, M is said to be Gorenstein flat if it can be written as M = Ker f for some long exact sequence

… F_{1} → F_{0} → F^{0} ⟶^{f} F^{1} → …

of flat modules which stays exact under I ⊗_{R} − for any injective module I_{R}. We denote by GI, GP, and GF the class of all Gorenstein injective, Gorenstein projective, and Gorenstein flat modules, respectively.

If R is a Gorenstein ring, then we know from [19, 11.2.3 and 10.1.2] that (P, GI) is a complete cotorsion pair whose kernel P ∩ GI is just the class of all injective modules. Similarly, (GP, P) is a complete cotorsion pair whose kernel GP ∩ P is just the class of all projective modules, see [19, 11.5.10 and 10.2.3].

The results above now show that over a Gorenstein ring, (P, GI) coincides with the tilting cotorsion pair of finite type cogenerated by P^{<∞}. Let us determine the corresponding cotilting cotorsion pair of cofinite type.

Theorem 3.4 The following statements are equivalent for a noetherian ring R.

(1) R is Gorenstein.

(2) The Gorenstein injective modules in ModR and in RMod form a tilting class.

(3) The Gorenstein injective modules in ModR form a tilting class and the Gorenstein flat modules in ModR form a cotilting class.

Proof : (1)⇒(2),(3): We know by Theorem 3.2 that GI = (P^{<∞})^{⊥} is a tilting class of finite type in ModR. Under the bijection of Theorem 2.2, the class GI is mapped to the cotilting class (P^{<∞})^{⊺} in RMod, which by Lemma 1.4.1 consists of the left modules M such that M^{∗} belongs to GI. By [19, 10.3.8] the latter are precisely the Gorenstein flat left modules. The statements on the Gorenstein flat right modules and on the Gorenstein injective left modules hold by symmetry. (2),(3)⇒(1): If GF is a cotilting class in ModR, then by [2, Lemma 2.2(b)] there is an integer n ≥ 0 such that every module X_{R} has a resolution 0 → F_{n} → … → F_{1} → F_{0} → X → 0 with Gorenstein flat modules F_{0}, …, F_{n}. We show that this implies idimR_{R} ≤ n. Recall that idimR_{R} is the supremum of the flat dimensions attained on the injective left modules, [19, 9.1.6]. So, let _{R}E be an injective module. By [19, 10.3.1] all Gorenstein flat right modules belong to ^{⊺}E, thus by dimension shift we have Tor_{i+n}^{R}(X, E) ≅Tor_{i}^{R}(F_{n}, E) = 0 for any module X and any i ≥ 0, and the claim is proven. Assume now that GI is a tilting class in ModR. Then by [2, Lemma 2.2] there is an integer n ≥ 0 such that ^{⊥}GI ⊆ P_{n}. Now, let E_{R} be an injective module. By [19, 10.1.3] all Gorenstein injective modules belong to E^{⊥}, hence we infer that E ∈ ^{⊥}(E^{⊥}) ⊆ ^{⊥}GI has (projective and) flat dimension at most n. As above, we conclude that idim_{R}R ≤ n. By symmetry, we finally obtain that idimR_{R} < ∞ provided that the Gorenstein injective left modules form a tilting class. □

Enochs and Jenda have proven in [20] that the Gorenstein injective modules over a Gorenstein ring can be characterized as the modules X such that Ext_{R}^{i}(L, X) = 0 for all i > 0 and all countably generated modules L of finite projective dimension. Our results above improve this characterization.

Corollary 3.5 Let R be a Gorenstein ring, and let X_{R} be a right R-module.

(1) X is Gorenstein injective if and only if Ext_{R}^{i}(L, X) = 0 for all i > 0 and all finitely generated modules L_{R} of finite projective dimension.

(2) X is Gorenstein flat if and only if Tor_{i}^{R}(X, L) = 0 for all i > 0 and all finitely generated modules _{R}L of finite projective dimension.

It is well known that over a Gorenstein ring, GP = GF if and only if R is artinian. We now characterize the artinian case in terms of cotilting modules.

Proposition 3.6 The following statements are equivalent for a Gorenstein ring R.

(1) R is right (or left) artinian.

(2) (GP, P) is a cotilting cotorsion pair.

(3) R_{R} is a cotilting module and ^{⊥}R_{R} = GP.

Proof : Let R be n-Gorenstein. Since (GP, P) is a complete cotorsion pair with P = I = I_{n}, we know from [2, Theorem 4.2] that (GP, P) is a cotilting cotorsion pair iff its kernel GP ∩P is closed under products. In other words, (2) just means that the class of the projective right modules is closed under products. By a theorem of Chase [14], the latter is equivalent to the fact that R is right perfect (and left coherent), which means that R is right (or left) artinian. So, we have shown (1)⇔(2). We now show (1)⇒(3): First of all, (1) implies that R is Σ-pure-injective, and so we even have ProdR =AddR. Then for any injective cogenerator W_{R} the projective resolution 0 → P_{n} → … → P_{1} → P_{0} → W → 0 yields an exact sequence as in (C3). Furthermore, all products of copies of R are projective, which proves (C2). Finally, we know that idimR_{R} = n, yielding (C1). So R_{R} is a cotilting module. Moreover, ⊥R =^{⊥}AddR, and the latter class coincides with ^{⊥}P = GP.□

Note that in general for a Gorenstein ring R, the class ^{⊥}R need not coincide with the class of Gorenstein flat modules, and R need not be a cotilting module.

Example 3.7 (1) Let R = ℤ. Then R^{IN} is flat, hence Gorenstein flat, but it does not belong to ^{⊥}R which is the class of Whitehead groups [21, Chapter XII, 1.3]. Moreover, GP = P_{0}. So, the question whether GP = ^{⊥}R is the Whitehead problem the solution of which depends on set theoretic assumptions, see [21, Chapter XIII] .

(2) Let p be a prime number and R = 𝕁_{p} the ring of all p-adic integers. Then^{⊥}R = GF is the class of all torsion-free 𝕁_{p}-modules, and it is a cotilting class which properly contains CogenR, see [5, 2.6] and [4, 2.4]. Thus R_{R} is not a cotilting module. □

Let us now turn to artin algebras.

Corollary 3.8 An artin algebra R is Gorenstein if and only if there is a cotilting module C ∈ modR such that GP = ^{⊥}C.

Proof : The only-if part follows from Proposition 3.6. For the if-part, we use [9, Theorem 5.5] stating that under the given assumption every X ∈ modR admits a finite GP ∩ modR-resolution. By [19, 10.2.6], the class GP ∩ modR consists of the reflexive modules G_{R} such that G ∈ ^{⊥}R_{R} and Hom_{R}(G, R) ∈ ^{⊥}(_{R}R), in other words, of the modules of Gorenstein dimension zero. So, we conclude that every finitely presented right module has finite Gorenstein dimension, which by [10, 24] means that R is Gorenstein.□

Note that GP ∩ modR is always a resolving class which is contained in X = ^{⊥}R_{R} ∩ modR, see [19, 10.2.8 and 10.2.2]. Moreover, if R is an artin algebra, then GP is a precovering class by [26]. However, we don’t know whether X is contravariantly finite in modR unless R is Gorenstein, cf. Proposition 3.1.

Recall from [3] that a module M satisfying condition (T1) and (T2) in the definition of a tilting module is called a partial tilting module. Furthermore, a module N is a complement of M if N ⊕ M is a tilting module. Partial cotilting modules and their complements are defined dually.

Lemma 3.9 Let R be an artin algebra with duality D : modR → Rmod. The following statements are equivalent.

(1) idim_{R}R = n < ∞.

(2) D(_{R}R)^{⊥} is a tilting class in ModR.

(3) ^{⊥}(_{R}R) is a cotilting class in RMod.

Proof : (1)⇒(2): The injective right R-module W = D(_{R}R)_{R} has projective dimension at most n and obviously satisfies condition (T2). So, it is a finitely presented partial tilting module, and by [3, Corollary 2.2] it has a complement M such that M ⊕ W is a tilting module with (M ⊕ W)^{⊥} = W^{⊥}. (2)⇒(3): By Proposition 2.2 we know that D(_{R}R)^{⊺} = ^{⊥}D^{2}(_{R}R) = ^{⊥}(_{R}R) is a cotilting class. (3)⇒(1): Of course _{R}R ∈ (^{⊥} _{R}R)^{⊥}, and by [2, Lemma 2.2] there is an integer n ≥ 0 such that (^{⊥} _{R}R)^{⊥} ⊆ I_{n}. □

Proposition 3.10 Assume that R is an artin algebra such that_{R}R has finite injective dimension. The following statements are equivalent.

(1) R is Gorenstein.

(2) D(_{R}R)^{⊥} = (P^{<∞})^{⊥}.

(3) The partial tilting module D(_{R}R) has a complement of finite injective dimension.

(4) The partial cotilting module _{R}R has a complement of finite projective dimension.

(5) Every (finitely presented) module in D(_{R}R)^{⊥} is generated by D(_{R}R).

(6) Every (finitely presented) module in ^{⊥}(_{R}R) is cogenerated by _{R}R.

(7) GP = ^{⊥}(_{R}R).

Proof : From Theorem 2.2 we know that (1) implies (2). Moreover, (1) implies (3) and (5) since D(_{R}R) is a tilting module and hence satisfies D(_{R}R)^{⊥} ⊆ GenD(_{R}R) by [2, Lemma 2.3]. Dually one obtains that (1) implies (4) and (6). Finally, (1) implies (7) by Proposition 3.6. For the converse implications, first of all observe that (2) implies that findimR is finite by Lemma 3.9 and [6, Theorem 2.6]. Then [9, 6.10] yields that idimR_{R} is finite. Next, assume (3). Condition (T3) then implies that R_{R} has finite injective dimension. Further, (7)⇒(5) since by definition GP ⊆ Cogen_{R}R, and (5)⇒(3) because the partial tilting module D(_{R}R) is then even tilting by [23, Proposition in §3] or [2, Theorem 4.4]. The remaining implications are proven dually. □

4 Commutative 1-Gorenstein rings

In this section, we first consider in greater detail a construction of tilting modules of finite type over commutative 1-Gorenstein rings:

Example 4.1 Let R be a commutative 1-Gorenstein ring. Let P_{i} denote the set of all prime ideals in R of height i (i = 0, 1). By a classical result of Bass, R has a minimal injective resolution of the form

0 ⟶ R ⟶ Q ⟶ ⨁_{p∈P_{1}} E(R/p) ⟶ 0

where Q = ⨁_{p∈P_{0}} E(R/p).

Recall from Corollary 3.3 that Q ⊕ ⨁_{p∈P_{1}} E(R/p) is the tilting module inducing the class of Gorenstein injectives. We are now going to consider a family of tilting modules of finite type obtained as variations of that construction, which in some cases lead to a classification of all tilting classes of finite type, cf. Remark 4.2.

Let P ⊆ P_{1}. Let F_{P} = ⨁_{p∈P} E(R/p) and G_{P} = ⨁_{q∈P_{1}\P} E(R/q). Consider the module R_{P} defined by R ⊆ R_{P} ⊆ Q and R_{P} /R ≅ F_{P}. We have the exact sequence 0 → R_{P} → Q → G_{P} → 0. Since the module G_{P} has flat dimension ≤ 1 and Q is flat (cf. [19, 9.1.10 and 9.3.3]), R_{P} is also flat. Define

T_{P} = R_{P} ⊕ F_{P} .

We will prove that T_{P} is a 1-tilting module of finite type. Clearly, T_{P} has projective dimension ≤ 1. We verify condition (T2) for T_{P}: since

R_{P} is {R} ∪ {E(R/p) | p ∈ P }-filtered, it suffices to prove that

Ext_{R}^{1} (E(R/p), R_{P}^{(κ)}) = 0

for all p ∈ P and any κ. Consider the exact sequence

0 ⟶ R_{P}^{(κ)} ⟶ Q^{(κ)} ⟶ G_{P}^{(κ)} ⟶ 0.

Applying Hom_{R}(E(R/p), −), we get part of the induced long exact sequence

0 = Hom_{R}(E(R/p), G_{P}^{(κ)}) → Ext_{R}^{1} (E(R/p), R_{P}^{(κ)}) → Ext_{R}^{1} (E(R/p), Q^{(κ)}) = 0

This proves (T2). The exact sequence 0 → R → R_{P} → F_{P} → 0 yields condition (T3) for T_{P}.

Altogether, T_{P} is 1-tilting; so Gen(R_{P}) = Gen(T_{P}) = T_{P}^{⊥_{1}} = F_{P}^{⊥_{1}} is a torsion class in ModR.

Let p ∈ P. Since R is 1-Gorenstein, Proposition 3.1 provides for a short exact

sequence 0 ⟶ R/p ⟶^{ρ} M_{p} ⟶ N_{p} ⟶ 0 such that M_{p} ∈ P_{1}^{<∞} and N_{p} ∈ ^{⊥_{1}}R.

Let (C_{n} | n < ω) be a composition series in E(R/p). By induction, we construct a direct system of short exact sequences 0 ⟶ C_{n} ⟶ D_{n} ⟶ E_{n} ⟶ 0 (n < ω) and connecting maps (μ_{n}, ν_{n}, π_{n}) such that μ_{n} is the embedding of C_{n} into C_{n+1} and ν_{n} is a monomorphism with Coker(ν_{n}) ≅ M_{p} for each n < ω.

For n = 0, we take the trivial exact sequence 0 → 0 → 0 → 0 → 0. Next, consider the push-out of the monomorphism C_{n} → D_{n} and of μ_{n}:

0  0
↓  ↓
0 ⟶ C_{n} ⟶ D_{n} ⟶ E_{n} ⟶ 0
μ_{n}↓  ^{α}↓  ∥
0 ⟶ C_{n+1} ⟶ F_{n} ⟶^{ρ} E_{n} ⟶ 0
↓  ^{β}↓
R/p  R/p
↓  ↓
0  0

Since D_{n} has projective (and injective) dimension ≤ 1, we have Ext_{R}^{2}(N_{p}, D_{n}) = 0. It follows that the map Ext_{R}^{1} (ρ, D_{n}) : Ext_{R}^{1} (M_{p}, D_{n}) → Ext_{R}^{1} (R/p, D_{n}) is surjective, so there is a pull-back diagram

0  0
↓  ↓
D_{n}  D_{n}
α↓  ↓
0 ⟶ F_{n} ⟶ D_{n+1} ⟶ N_{p} ⟶ 0
β↓  ↓  ∥
0 ⟶ R/p ⟶^{ρ} M_{p} ⟶ N_{p} ⟶ 0
↓  ↓
0  0

Altogether, we get a commutative diagram with exact rows and columns that provides for the induction step of the construction

0  0  0
↓  ↓  ↓
0 ⟶ C_{n} ⟶ D_{n} ⟶ E_{n} ⟶ 0
μ_{n}↓  ^{ν_{n}}↓  ^{π_{n}}↓
0 ⟶ C_{n+1} ⟶ D_{n+1} ⟶ E_{n+1} ⟶ 0
↓  ↓  ↓
0 ⟶ R/p ⟶^{ρ} M_{p} ⟶ N_{p} ⟶ 0
↓  ↓  ↓
0  0  0

Let 𝔇 : 0 ⟶ E(R/p) ⟶ D ⟶ E ⟶ 0 be the direct limit of the system 0 ⟶ C_{n} ⟶ D_{n} ⟶ E_{n} ⟶ 0 (n < ω). Since E(R/p) is injective, 𝔇 splits. Since Coker(ν_{n}) ≅ M_{p} for each n < ω, the module D is {M_{p}}-filtered. In particular, we have E(R/p)^{⊥_{1}} ⊇ M_{p}^{⊥_{1}}.

Let S_{P} = {M_{p} | p ∈ P }. We have just shown that T_{P}^{⊥_{1}} ⊇ S_{P}^{⊥_{1}}.

In order to prove that T_{P} is of finite type, it suffices to show that T_{P}^{⊥_{1}} ⊆ M_{p}^{⊥_{1}} for all p ∈ P. For this, it is enough to prove that Ext_{R}^{1} (M_{p}, R_{P}) = 0 for each p ∈ P (then Gen(R_{P}) ⊆ M_{p}^{⊥_{1}}). Applying Hom_{R}(−, R_{P}) to the exact sequence 0 → R/p → M_{p} → N_{p} → 0, we get

… ⟶ Ext_{R}^{1} (N_{p}, R_{P}) ⟶ Ext_{R}^{1} (M_{p}, R_{P}) ⟶ Ext_{R}^{1} (R/p, R_{P}) ⟶ …

Applying Hom_{R}(N_{p}, −) to the exact sequence 0 → R → R_{P} → F_{P} → 0 and using the fact that N_{p} ∈ ^{⊥_{1}}R, we get Ext_{R}^{1} (N_{p}, R_{P}) = 0. Since Hom_{R}(R/p, Q/R_{P}) = 0, we have Ext_{R}^{1} (R/p, R_{P}) = 0. It follows that Ext_{R}^{1} (M_{p}, R_{P}) = 0.

This proves that T_{P} = T_{P}^{⊥_{1}} = S_{P}^{⊥_{1}} is a 1-tilting class of finite type.

Applying Proposition 2.3, we can now describe the corresponding 1-cotilting modules, and 1-cotilting classes, of cofinite type.

We will use the duality (−)^{∗} = Hom_{R}(−, W) induced by the injective cogenerator W = ⨁_{q∈P_{0}∪P_{1}} E(R/q). For each P ⊆ P_{1}, we obtain a 1-cotilting module

C_{P} = T_{P}^{∗} = D_{P} ⊕ ∏_{p∈P} J_{p}

where D_{P} = R_{P}^{∗}, and J_{p} = Hom_{R}(E(R/p), E(R/p))) = R̂_{p} is the p-adic module. Since R_{P} is flat, the module D_{P} is injective.

We claim that D_{P} ≅ F_{P} ⊕ ⨁_{q∈P_{0}} E(R/q)^{(α_{q})} where α_{q} ≥ 1 for all q ∈ P_{0}. Indeed, D_{P} ≅ Hom_{R}(R_{P} , Q⊕G_{P})⊕Hom_{R}(R_{P} , F_{P}). But Hom_{R}(R_{P} , Q⊕G_{P}) ≅ Hom_{R}(R, Q⊕ G_{P}) ≅ Q ⊕ G_{P}, while Hom_{R}(R_{P} , F_{P}) ≅ Hom_{R}(Q, F_{P}). The latter module is flat and injective, hence isomorphic to a direct summand of Q^{(J)} for a set J, and the claim follows.

The corresponding 1-cotilting class is F_{P} = ^{⊥_{1}}C_{P} = S_{P}^{⊺_{1}} (cf. Theorem 2.2.3). □

Remark 4.2 (i) The construction of the module M_{p} in Example 4.1 can be made more explicit for any p ∈ P satisfying p ⊈ ∪_{q≠p}q (By prime avoidance, the latter holds for all p ∈ P in case P_{1} \ P is finite.) For such p there is a regular element y_{p} ∈ p \ ⋃_{p≠q∈P_{1}} q. Since R is 1-Gorenstein, y_{p} forms a maximal p-sequence in R and we can take M_{p} = R/y_{p}R. The proof is then the same as above except that Ext_{R}^{1} (M_{p}, R_{P}) = 0 follows from Hom_{R}(M_{p}, R/q) = 0 which in turn is a consequence of y_{p} being an R/q-regular element in y_{p}R.

(ii) The reasoning in Example 4.1 can substantially be simplified in case R is a Dedekind domain. Though in this case it can still happen that p ⊆ ∪_{q≠p}q (so part (i) may not apply), we have a much simpler choice: M_{p} = R/p. It follows that T_{P} is just the class of all modules that are p-divisible for each p ∈ P, and F_{P} is the class of all modules that are p-torsion-free for each p ∈ P. Moreover, the classes F_{P} (P ⊆ P_{1}) are the only cotilting classes of modules by [11] and [17, Corollary 17], and the classes T_{P} (P ⊆ P_{1}) are the only tilting classes of finite type, cf. [32, Theorem 5]. If R is a small Dedekind domain and V=L, then the classes T_{P} (P ⊆ P_{1}) are the only tilting classes of modules, [32, Theorem 12].

We finish this part by considering tilting modules and tilting classes of finite type under classical localization. Let R be a commutative ring, S ⊆ R a multiplicative set, T ∈ ModR, and T ⊆ ModR. Then R_{S} denotes the localization of R at S, T_{S} ≅ T ⊗_{R} R_{S} the localization of T at S, and T_{S} = {N ∈ ModR_{S} | N ≅ M ⊗_{R} R_{S} for some M ∈ T }. As usual, for a maximal ideal m of R, T_{m} = T_{R\m} and T_{m} = T_{R\m}.

Proposition 4.3 Let R be a commutative ring. Let T be an n-tilting module of finite type and T = T^{⊥} be the corresponding n-tilting class.

1. Let S be a multiplicative subset of R. Then T_{S} is an n-tilting R_{S}-module of finite type, the corresponding n-tilting class being T_{S} = T_{S}^{⊥} = T ∩ ModR_{S}.

2. Let M ∈ ModR. Then M ∈ T if and only if M_{m} ∈ T_{m} for all maximal ideals m of R.

Proof : By assumption, there is a resolving class C ⊆ P_{n}^{<∞} such that T^{⊥} = C^{⊥_{1}}. Let 𝔓 be a projective resolution of T in ModR. By [31], each syzygy of T in 𝔓 is

isomorphic to a direct summand of a C-filtered module, and conversely, each module in C is isomorphic to a direct summand of a module filtered by the syzygies of T.

1. Applying the exact functor − ⊗_{R} R_{S} to the long exact sequences in (T1) and (T3) for T, we obtain conditions (T1) and (T3) for T_{S}.

Further, each element of C_{S} has a projective resolution of length ≤ n consisting of finitely generated R_{S}-modules.

Localizing 𝔓 at S, we obtain a projective resolution 𝔔 = 𝔓⊗_{R}R_{S} of T_{S} in ModR_{S}. Since R_{S} is a flat module, each syzygy of T_{S} in 𝔔 is isomorphic to a direct summand of a C_{S}-filtered R_{S}-module, and conversely, each R_{S}-module in C_{S} is isomorphic to a direct summand of an R_{S}-module filtered by the syzygies of T_{S}. By [31], this just says that in ModR_{S}, T_{S}^{⊥} = C_{S}^{⊥_{1}}.

Let i ≥ 1 and let I be a set. Condition (T2) for T and (a generalization of) [19, 3.2.5] now give

0 = Ext_{R}^{1} (C, T^{(I)}) ⊗_{R} R_{S} ≅ Ext_{R_{S}}^{1} (C_{S}, T_{S}^{(I)})

for each C ∈ C. It follows that T_{S}^{(I)} ∈ T_{S}^{⊥}, so condition (T2) holds for the R_{S}- module T_{S}.

Similarly, [19, 3.2.5] gives T_{S} ⊆ C_{S}^{⊥_{1}} (= T_{S}^{⊥}), and since N ≅ N ⊗_{R} R_{S} for any module N ∈ ModR_{S} (⊆ModR), we also get N ∈ C_{S}^{⊥_{1}} iff N ∈ C^{⊥_{1}}. It follows that T_{S}^{⊥} = T ∩ ModR_{S}, and clearly T ∩ ModR_{S} ⊆ T_{S}.

Since C_{S} ⊆ modR_{S}, the equality T_{S}^{⊥} = C_{S}^{⊥_{1}} gives that T_{S} is of finite type.

2. By (a generalization of) [19, 3.2.15], we have

Ext_{R}^{1} (C, M) ⊗_{R} R_{m} ≅ Ext_{R}^{1} (C, M_{m})

for each C ∈ C. Since T = C^{⊥_{1}}, M ∈ T iff M_{m} ∈ C^{⊥_{1}} for all maximal ideals m of R. The latter is equivalent to M_{m} ∈ C_{m}^{⊥_{1}} = T_{m} for all maximal ideals m of R. □

References

[1] K. Ajitabh, S. P. Smith, J. J. Zhang, Auslander-Gorenstein rings, Comm. Algebra 26 (1998), 2159–2180.

[2] L. Angeleri Hügel and F.U. Coelho, Infinitely generated tilting modules of finite projective dimension, Forum Math 13 (2001), 239-250.

[3] L. Angeleri Hügel and F.U. Coelho, Infinitely generated complements to partial tilting modules. Math. Proc. Cambridge Philos. Soc. 132 (2002), no. 1, 89–96.

[4] L. Angeleri Hügel, Finitely cotilting modules, Communications in Algebra 28 (2000), 2147-2172.

[5] L. Angeleri Hügel, A. Tonolo and J. Trlifaj, Tilting preenvelopes and cotilting precovers, Algebras and Repres. Theory 4 (2001), 155-170.

[6] L. Angeleri Hügel and J. Trlifaj, Tilting theory and the finitistic dimension conjectures, Trans. Amer. Math. Soc. 354(2002), 4345-4358.

[7] L. Angeleri Hügel and J. Trlifaj, Direct limits of modules of finite projective dimension, to appear in Proc. Algebra Conf. Venezia 2002, LNPAM, M. Dekker.

[8] M. Auslander, R. O. Buchweitz, The homological theory of maximal Cohen-Macaulay approximations, Memoires of Soc. Math. de France (Suppl.), Nouvelle Serie 38 (1989) 5-37.

[9] M. Auslander, I. Reiten, Applications of contravariantly finite subcategories. Adv. Math. 86 (1991), 111-152.

[10] M. Auslander, I. Reiten, Cohen-Macaulay and Gorenstein Artin algebras, Progress in Mathematics 95 (1991), 221-245.

[11] S. Bazzoni, Cotilting modules are pure-injective, to appear in Proc. Amer. Math. Soc.

[12] K.S. Brown, Cohomology of Groups, Springer, New York 1982.

[13] W. Bruns and J. Herzog, Cohen-Macaulay Rings, 2nd ed., CSAM 39, Cambridge, 1998.

[14] C. U. Chase, Direct product of modules, Trans. Amer. Math. Soc. 97 (1960), 457-473.

[15] W. Crawley-Boevey, Locally finitely presented additive categories, Comm. Algebra 22 (1994), 1644-1674.

[16] P. Eklof, J. Trlifaj, How to make Ext vanish, Bull. London Math. Soc. 33 (2001), 41-51.

[17] P. Eklof, J. Trlifaj, Covers induced by Ext, J. Algebra 231 (2000), 640-651.

[18] E. Enochs and O. Jenda, Gorenstein injective and projective modules, Math. Zeitschrift 220 (1995), 611-633.

[19] E. Enochs and O. Jenda, Relative Homological Algebra, de Gruyter, Berlin, 2000.

[20] E. Enochs and O. Jenda, Gorenstein injective modules and Ext, preprint.

[21] P. C. Eklof and A. H. Mekler, Almost Free Modules: set-theoretic methods, North-Holland (1990).

[22] L. Fuchs and L. Salce, Modules over Non-Noetherian Domains, AMS (2001).

[23] D. Happel, Selforthogonal modules. In: Abelian Groups and Modules. Proc. of Padova Conf. (ed. by A.Facchini et al.) Math. Appl. bf 343 (1995), 257-276.

[24] M. Hoshino, Algebras of finite self-injective dimension, Proc.Amer.Math.Soc. 112 (1991), 619-622.

[25] K. Igusa, S.O.Smalø, G. Todorov, Finite projectivity and contravariant finiteness, Proc. Amer. Math. Soc. 109 (1990), 937-941.

[26] P. Jørgensen, Spectra of modules, J. Algebra 244, (2001) 744-784.

[27] H. Lenzing, Homological transfer from finitely presented to infinite modules, Lecture Notes in Mathematics 1006, Springer, New York 1983, 734-761.

[28] J. Miyachi, Duality for derived categories and cotilting bimodules, J. Algebra 185 (1996), 583-603.

[29] J. Miyachi, Injective resolutions of noetherian rings and cogenerators, Proc. Amer. Math. Soc. 128 (2000), 2233-2242.

[30] Y. Miyashita, Tilting modules of finite projective dimension, Math. Zeitschrift 193 (1986), 113-146.

[31] J. Trlifaj, Cotorsion theories induced by tilting and cotilting modules, Contemp. Math. 273 (2001), 285-300.

[32] J. Trlifaj and S. Wallutis, Tilting modules over small Dedekind domains, J. Pure. Appl. Algebra 172 (2002), 109-117.

Address of the authors:

Lidia Angeleri Hügel  Dolors Herbera  Jan Trlifaj

Dipartimento di Informatica Departament de Mathemàtiques Katedra algebry MFF

e Comunicazione, Università  Universitat Autonoma  Univerzita Karlova
degli Studi dell’Insubria  de Barcelona  v Praze
Via Mazzini 5, I - 21100 Varese  08193 Bellaterra  Sokolovská 83, 186 75 Prague 8
Italy  Spain  Czech Republic
e-mail: lidia.angeleri@uninsubria.it  dolors@mat.uab.es  trlifaj@karlin.mff.cuni.cz